Oxford Materials IT Committee Meeting

TRINITY TERM 2006
A meeting of the Department of Materials IT Committee was held on Tuesday 9th May 2006.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters arising from the Minutes of the Meeting of 31st January 2006
  • 3 IT/AV – Matters for Discussion (attached)
  • 4 IT/AV – Head of IT Report (attached)
  • 5 University IT Users Group – brief verbal report (PJW)
  • 6 IT/AV expenditure - matters for report (PJW)
  • 7 AOB
  • 8 Date and time of next meeting: Tuesday 24th October 2006 at 2pm.

ITEM 2. MATTERS ARISING FROM THE MEETING

See minutes below.

ITEM 3: IT/AV MATTERS FOR DISCUSSION

Oxford Wireless Lan - discussion of prelimary proposals for implementing OWL in department buildings.

ITEM 4: IT/AV HEAD OF IT REPORT

Admin Network
Now-upto-date shared calendar software installed for APD+AJJ. Peter Flaxman, GDWS and ICS's PC's were upgraded by installing a new PC for ICS and then trickle-down to GDWS and PF. This also removed the last WindowsNT workstation (GDWS) so that all the Admin Network is now running WindowsXP.

Teaching Laboratory
Four new computers (Dell 5150 P4 3.2GHz) were installed into the teaching laboratory computer room and the old computers (Dell GX240 P4 1.5GHz) were trickled to replaced even older computers elsewhere (Librarian, LibrarySelfIssue, XRD, and BegbrokeWriteUpPC).

Lecture theatres / meeting rooms
The CETL equipment (smartboard, projector, laptop, trolley, two slates and a sympodium touchsensitive monitor) was collected from Physics. The sympodium was installed in the Hume-Rothery Lecture Theatre. The Smartboard and other associated equipment was installed in ETB Meeting Room (formerly LR7). Training of staff to use these new teaching aids is being organised by Physics for later this term as part of the CETL project.
Plans for the new ITAV equipment as part of the refurbishment of 21 Banbury Road Lecture Theatre over the summer are being finalized and quotes obtained ready for a funding meeting with the UED. [ dual dataprojectors, vizualizer, sympodium, possible video link to Begbroke, etc.]

Digital Darkroom
Intermittent smearing problems with the HP8550 colour printer continue to be fixed by Andrew McKnight (once per month).

Library
The self-issue software interface was upgraded prior to the major changes in OLIS scheduled for this summer. The Self-Issue computer and the Librarian's computer were both upgraded by tricke-down from the Teaching Laboratory.

MML
The MML working area was re-organised after the department's annual computer equipment disposal in March. The test cluster was dismantled and all the computers returned to the normal queuing system. The migration of all workstations to SuSE10 was completed. The 48-node cluster was specified and purchased costing £25k. The cluster was installed in the Hume-Rothery Plantroom in March but continues to have hardware instabilities which mean that the system has not yet been made available to MML users. Viglen are due to attend site again soon. This has been very time-consuming.

EM
New SGI's were purchased and reinstalled/reconfigured by KS. A method for access to the Department-wide Altiris system for backup of EM PC's is still being investigated. The websem webserver fried but was replaced under warrantee and the software recovered and a new server purchased for development. The emgroup1 server is due for decommissioning as soon as users/data can be transferred to Materials0. The EM firewall was updated/replaced.

Website
The intended review of research group websites last summer still has not happened. PJW and JAMS met and discussed the requirements for improving research group websites. Several Departmental newsfeeds were created making use of the new newsfeed system implemented by OUCS.

Publications
Multiple ~12 copies of all department publications were produced in-house for the EPSC departmental review. Invitation cards (~2000 copies) for the 50th Anniversary event in October were produced out-of-house. As usual requests were dealt with efficiently despite short notice! Business cards continue to be requested frequently.

Network
The new department firewall seems to have been working well and a methodology for daily log-checking and refinement of rules is continuing. A switch hardware failure in HR was replaced within 1hr. A weekend power-spike at Begbroke caused a switch to require a manual restart after the weekend. Switches were upgraded in the HR main stack and at Begbroke HirschSouthExtension. The FROntDOor project to install active devices in all buildings under the control of OUCS has slipped to the summer. The re-wiring in the basement of ETB is complete but the network feeds in ETB still need to be re-routed. Re-wiring of 21BR is half-way through the planned works.

Servers
Materials2(main netware fileserver) had several large software upgrades affecting all clients/users.
Oums-nomad (Altiris Deployment Server, antivirus CID, FM Group database, NUTD) now also serves Now Up-to-date.
Oums-wsus (windowsupdates) server has been installed and is working well.
Resolve2 has been decommissioned and is being reused for network monitoring.
Databases(the old sophos) server now only has a few (6) people connecting to it.
Oums-sus(the old windowsupdates) server now only has a few (7) people connecting to it.
www-int, www-ext (webservers) no changes.
Test installations using Novell OES Linux and Novell OES Netware are being run on Materials1 and ittemp3. Materials0 was reinstalled, the disk capacity expanded to 3Tb and is now ready for user-testing.

Phones
There have been very few reports of problems with the phones. Most common problem is drop-out of one or both audio signals during a conversation, which is solved by calling back (however this is clearly inconvenient especially when dealing with call centres). Problems reported to Telecoms.

Security
No computers were hacked. One user was reprimanded as a repeat offender. Five other instances of possible misuse were investigated but not confirmed.


MINUTES

A meeting of the Department of Materials IT Committee was held on Tuesday 9th May 2006.
The following people attended: APD, PJW, KS, JAMS, CJS.

1. Apologies for Absence and Membership of the Committee

Apologies from AHM - missed due to overlapping meetings.

2. Matters arising from the minutes

a) Email and shared calendar server (Minute 2b)
Requirements of an Exchange server have been investigated. Informal comments on Lotus Notes as used in Central Admin were unfavourable. Other systems were considered but all had pros and cons with there being no clear optimum system. Groupwise remains of interest but current needs for shared calendaring do not justify the time necessary to implement it.
Our recommendation to users is to make use of the new pay-per-use Exchange server system run by NSMS(OUCS). Department IT staff will review this advice as needs arise.
A free-standing shared calendar system (Now Up-to-date) has been implemented for APD+AJJ since APD used this system previously in Chemistry. This system can be offered to other users subject to paying for additional licenses.
DJHC+KH enquired about shared calendaring for Outlook and were directed to the new pay-per-use Exchange server system run by NSMS(OUCS) but they have opted to use a windows file-share to achieve a shared calendar.
CJS mentioned that the OMCS team at Begbroke are actively looking for a calendaring solution. PJW/AMC/CJS to discuss.

Action: PJW/AMC/CJS

b) IT disaster recovery plans (Minute 2c)
PJW circulated, with the agenda, a draft review of UPS provision within the department as the first step towards creating disaster recovery documentation.
The department has spent several thousands of pounds over the past year improving UPS cover for servers and network switches but further work is needed in this area so the item is still carried to the next meeting.

Action: PJW

c) Website (Minute 2f)
The proposed review of research group websites did not happen during the summer but has been carried forward to the next meeting. Advice and guidance is to be offered by PJW upon compliance with University Website Policies and upon mechanisms for easy updating of content. PJW to email notices to raise interest in web authoring and best practice.

Action: PJW

d) Copyright (Minute 7a)
APD noted that new copyright regulations need to be assessed and staff notified of changes. PJW and Adrian Taylor discussed the new copyright regulations. OUCS are running a course 'WISER Teaching: Copyright Law' on 7th week. Tuesday 6 June 12:30-13:30. JAMS asked PJW to check whether it was relevant to senior staff and also to find out what compliance procedures and what enforcement procedures exist.

Action: PJW

f) Website (Minute 2f)
JAMS raised several points concerning the department website. The new projects page needs to be clearer for prospective students to find projects of interest. The current status of representation of research groups on the Department website is patchy at best, and a centralized system for summarizing Research Group information is desirable. PJW held a first meeting with JAMS and is currently drafting a template for a research group website to be rolled out over the summer.

Action: PJW/JAMS

Ian Atkin queried why all the IT section of the website was restricted to only being visible from within Materials. PJW assessed which content (e.g. policies) should be more widely accessible and decided that all IT information should be available to all Oxford, but none available to the World.

3 Matters for Discussion

The implication of the new Oxford Wireless LAN was discussed.
As a result of the FRODO project putting active equipment under the control of OUCS inside the FROntDOor is is now possible to make the Oxford Wireless LAN available within departmental buildings. This will involve some technical reconfiguration of our existing network devices and the addition of Wireless Access Points at suitable locations in order to give coverage.
A suggested project to implement Oxford Wireless LAN over the summer was greeted enthusiastically. The first stage will be to provide coverage for meeting rooms, lecture theatres, coffee areas and replacing the pilot wireless project in the QIPIRC. Implicitly this will require on-the-job training for department IT staff. The second stage will be to negotiate with group leaders about provision in their respective group areas and what funds are available for hardware and installation. There will be need for co-ordination with Engineering and the Begbroke Directorate over coverage in shared buildings.

4. Head of IT Report

A written report was circulated with the agenda (above).
APD noted that the new procedures for dealing with computer misuse by students (formulated by AOT in response to a repeat offence of p2p filesharing) need to be presented to DC on Monday 12th June.

Action: PJW/AOT

5. Oxford University IT Users Group Meeting http://www.ict.ox.ac.uk/oxford/groups/ouitug/

No specific items for report.

It was noted that at the next meeting it will be proposed that the ITUG be disbanded and the role and responsibilities be taken by the new ICT Forum as proposed in the new ICT strategy document currently awaiting approval. http://www.ict.ox.ac.uk/strategy/.

6. IT/AV Expenditure

PJW apologised for not having finished preparing a status report on IT expenditure ready for the meeting.
A summary of the expenditure for 2006 was later circulated with the minutes of the meeting. The total expenditure for the financial year 1st August 2005-31st July 2006 is likely to be around £40k. A breakdown of expenditure in OSIRIS to-date by category was also circulated. Expenditure for next financial year 2006/7 is forecast to be similarly around £40k but detailed financial implications of the proposed wireless network (Minute 3) are not available yet.

7 AOB

a) Microwave
APD asked for an update on the status of the microwave equipment. The equipment has been turned off and the license discontinued but there has been no progress in finding a buyer. PJW should sell or dispose of the equipment before the next meeting.

Action: PJW/KS

b) APL
APD pointed out that the new APL building at Begbroke is due to be officially completed this month. PJW confirmed that he had previously discussed networking requirements with PSG. PJW and KS to install a simple UTP network feed from Highbay Annex in order for Robin Vincent to have computer and VOIP phone links to central site.

Action: PJW/KS

8 Date of Next Meeting

The next meeting will take place on Tuesday 24th October at 2pm in 21BR Committee Room (room booked already)

The meeting closed at 3.20pm